One number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers.
Answer by checkley71(8403)

X=Y+6
Y+3X=34 SUBSTITUTING (Y+6) FOR X IN THIS EQUATION THEN SOLVE FOR Y
Y+3(Y+6)=34
Y+3Y+18=34
4Y=34-18
4Y=16
Y=16/4
Y=4 ANSWER.
X=4+6
X=10 ANSWER.
PROOF
4+3*10=34
34=34